OpenStudy (anonymous):

When you see a traffic light turn red, you apply the brakes until you come to a stop. If your initial speed was 10 m/s, and you were heading due south, what was your average velocity during braking? Assume constant deceleration.

OpenStudy (shane_b):

Your average velocity, given constant (+/-) acceleration is:\[V_{avg}\frac{V_{initial}+V_{final}}{2}\]Plugging in what you were given:\[V_{avg}\frac{10m/s+0m/s}{2}=5m/s\]